OpenStudy (anonymous):

1.4+2/5e=3/15e-0.8 Can someone help me work this out?

OpenStudy (coconutjj):

\[1.4+\frac{ 2 }{ 5 }e = \frac{ 3 }{ 15 }e - 0.8\] is this the equation

OpenStudy (anonymous):

Yes and e=-11 if that helps

OpenStudy (coconutjj):

yes.. haven't you already solved this problem

OpenStudy (anonymous):

No the teacher gave the answers and said work it out.

OpenStudy (anonymous):

So, I need help working this one out I finished all the rest but this one has me stomped.

OpenStudy (coconutjj):

hmmm... what's troubling you

OpenStudy (anonymous):

The fractions so do I start by subtracting 1.4 from both sides?

OpenStudy (coconutjj):

YES.. move 1.8 to the other side... so you get ...=(3/15) -0.8 -1.4

OpenStudy (anonymous):

I get -0.6

OpenStudy (coconutjj):

you should get -2.2... (-0.8 - 1.4) = ?

OpenStudy (anonymous):

Yes -2.2 ok got it

OpenStudy (coconutjj):

\[\frac{ 2 }{ 5 }e = \frac{ 3 }{ 15 }e -\frac{ 11 }{ 5 }\] should be something like this

OpenStudy (anonymous):

Subtracted wrong I see now

OpenStudy (anonymous):

Ok

OpenStudy (coconutjj):

do you know how to do it from here

OpenStudy (anonymous):

Now do I combine the terms?

OpenStudy (coconutjj):

yes

OpenStudy (coconutjj):

2/5e - 3/15e = ??

OpenStudy (coconutjj):

ummm.... 2/5 - 3/15 = 1/5

OpenStudy (anonymous):

Omg I knew that I'm sleep deprived.

OpenStudy (coconutjj):

let hope you know how to solve it from here now... shall we?

OpenStudy (anonymous):

I will divide -2.2 by 0.2 and get -11

OpenStudy (anonymous):

Thank you